UT Sep 03 2019 - Restart Night#2


OSA: D.G. Huerta

ISp: G. Bechetti

Mtn Manager: P. Hartley

ISA: B. Rothberg, O. Kuhn (Tucson)
AO: D. Miller (Tucson)

ACCOMPLISHED:
  • LBCs
    • LBC Collimation
    • LBC Rotator collection and verification (old values OK)
    • Test Science Sidereal and Non-Sidereal
  • MODS
    • AGW-6 rotator center verification (data collected - need to confirm hotspot is correct)
    • MODS-2 sky flats (u,g,r,z)
    • MODS-2 dual and red grating spectro-std
    • MODS-2B operating without electronic readout issues

Rollover of tasks attempted last night
  • Telescope Wake-up ~0.5hr (test presets, TCS/PCS testing)
  • LBC Wake-up- contingent on LBCs able to collimate
    • DONE Collimation - DoHybrid, then DOFPIA (~1hr)
    • Superfocus Sequence, PA=0, 180 (1hr)
    • DONE Rotator Center Checks (pairs of 0,180; 90, 270) ~0.5hr
    • Test Science OB ~ 1hr
      • DONE Sidereal
      • DONE Non-Sidereal
    • Photometric Fields and Spectrophotometric Stars (to establish ZPs) ~ 1-1.5hrs
      • Requires multiple airmass visits (clear conditions, any seeing)
      • Cycle through all filters
      • This is now a high priority because we need to define ZPs
    • Twilight Flats
  • MODS On-sky checkout
    • DONE Confirm centering of AGWs and guiding (test preset)
    • Basic Checkout
      • Rotator Centers
        • Collimate and then send track presets to the same field at PA=0, 180 and PA=90, 270. Analyze data with modsrc . modsrc works on both MODS and LBC images and it computes rotator centers for pairs of images which do not have to be taken at diametrically opposed position angles but can be, say, 90 deg apart instead.
      • Rotator Zero Points
        • Acquire 2 MOS masks and note the delta theta offset needed to align the mask. Ideally, the rotator zero point will be set (in PCSInstrument.conf) so that the mask or longslit is oriented at PA=0. The delta theta by which the instrument needs to be rotated to align the mask and field is the error in the rotator zeropoint. If the rotator zeropoint is updated, then a new transform needs to be determined. A 0.7 deg error in the rotator zeropoint translates to 0.73" at a distance 60" from the slit center. (If one of two stars separated by 60" is perfectly centered in a long-slit slit, the other will be on one edge of the slit unless the field is rotated by 0.7 degrees; since MOS mask alignment computes a rotational offset as well as a shift, the error in teh rotator zeropoint is corrected during a MOS alignment).
      • Transform: At minimum, check the existing transform by making a series of DETXY offsets with large (60") throw to compare orientation of deltaX, deltaY with slit.
      • DONE (but partial only MODS-2 Dual and Red) Observe a standard star in all modes (30-45 min for 3 images). If clear, obtain the full set of 3 images per mode; if not, make a note in the object name and obtain only 1 image per mode.
      • Dice-5 imaging offsets
      • low priority check on orientation and scale of DETXY offsets, repeatability of these and also an opportunity to gather data on image quality.
    • Test Science OB
      • Sidereal
      • Non-Sidereal
  • AGW1 and AGW2 checkout and alignment at LFBG (critical)
    • initial wake-up
    • wfs alignment (adjust guiding hotspot in twilight, measure pupil position)
    • align probe to put hotspot on rotator center
    • adjust M3 to minimize pupil wobble (critical)
    • refine wfs alignment
  • measure LFBG/RFBG field aberrations (high)

UT 01:04 - blue camera failed to come up on LBC-B

camera system encountered a problem: CCD Controller error (9) [src/channel/channel.c:527]

-turning off LBCs and starting again

UT 01:10 - LBCs on

UT 01:11 - taking test Biases

- both bias -looks like too few pixels - taking more biases

UT 01:16 - running up MODS

UT 01:20 - last of the 2nd set of LBC-B biases look OK

UT 01:23 - LUCI-1 released for engineering (flexure and wavelength still TBD)

UT 01:52 - error co-pointing limits (and yes we are closed and authorized as LBCs)

UT 01:59 - we were about to open but a cloud appeared overhead so aborting that

UT 02:09 - sievesnap taken for both MODS, MODS-2B looks OK

UT 02:10 - taking maskSnap for three MOS masks

UT 02:22 - maskSnaps look good - MODS2B still OK

-going to do IMCS Zeropoints for both MODS since we are closed

-moving MODS-2 to 195, MODS-1 at 210, at elevation = 60

UT 02:45 - IMCS TT Zp completed, O. Kuhn now comparing new values with current config

UT 03:34 - MODS GUIs restarted after IMCS values updated by O.Kuhn

UT 03:35 - proceeding with MODS Cals

UT 04:22 - still closed

UT 04:41 - Opening Enclosure - moved while taking MODS calibrations data - arc lamps (comps) - kind of a surprise - no warning that we opened

UT 04:44 - MODS-1 script hung during cal lamps

UT 04:48 - going to RA 20:30:01 +59:59:29.9 for collimation

UT 04:51 - running dohybrid for collimation

UT 04:56 - got error that couldn't find data in /newdata

-LBC-B thinks its astigmatic - appears dohybrid is selecting the bright spot and assuming the hole in the pupil is in there. (The WRS model is a very elongated image, which matches the bright area of the comatic pupil. WRS has therefore added a lot of astigmatism which FPIA is left to take out).

-LBC-R - looks better --> however DX has +2000 z11 in global offsets - we didn't see this last night

-LBC-B is elongated, a lot of coma and astigmatism

UT 05:05 - dohybrid is done

UT 05:09 - copointing.

-We notice that PSF-R has a z11 global offset of 2000 and has from the start of the night, even though we authorized LBCs from scratch. During the night last night, Barry added z11 of 2000 to DX, but he also added some other aberrations as we were trying to achieve collimation. The others were cleared, but not z11.

UT 05:11 - running dofpia - image for co-pointing was fuzzy
error again:

******* WARNING *******
Data files DID NOT appear in /newdata
Pls double check and then: press Q to exit or any other key to continue
:
-however, the data were in /newdata

UT 05:17- LBC-R reported too many aberrations and unable to compute sky

-says it converged but the central hole in pupil is quite pinched

UT 05:19 - doing the co-pointing despite the less than optimal collimation

UT 05:24 - co-pointing:

Old Delta New
------- ------- -------
CE -37.608 -3.793 -41.401
CA -109.506 -0.313 -109.818
On LBC-Blue/LBC-Red the star is 9.8/15.6 arcsec away from the rotator center.Based on these, the following mount and optics offsets were calculated to co-point while optimizing the range available to M1:
MOUNT: delta_IE = -3.79 arcsec, delta_CA = -0.31 arcsec
M1 SX: dX = -0.27 mm, dY = 0.54 mm, dRX = 11.52 arcsec, dRY = 5.73 arcsec
M1 DX: dX = 0.25 mm, dY = -0.51 mm, dRX = -10.89 arcsec, dRY = -5.47 arcsec

Do you want to send these offsets? Answer 'y' or 'Y' if so. Any other key will abort.
y
Sending the calculated optics offsets:
SX: dX = -0.267 dY = 0.537 dRX = 11.520 dRY = 5.730
DX: dX = 0.255 dY = -0.507 dRX = -10.887 dRY = -5.468
Input term name: IE Input term value: -41.4007

SUCCESS

*** Pointing Model Terms **
IA -331.10000
IE -41.40074
NPAE 46.94600
CA -109.50570
AW 6.77230
AN 20.51350
PZZ1 32.89200
IA -331.10000
IE -41.40074
NPAE 46.94600
CA -109.50570
AW 6.77230
AN 20.51350
PZZ1 32.89200

Input term name: CA Input term value: -109.818

SUCCESS

**
Pointing Model Terms ***
IA -331.10000
IE -41.40074
NPAE 46.94600
CA -109.81830
AW 6.77230
AN 20.51350
PZZ1 32.89200
IA -331.10000
IE -41.40074
NPAE 46.94600
CA -109.81830
AW 6.77230
AN 20.51350
PZZ1 32.89200

UT 05:25 - Taking Rotator Center data lbcb.20190903.;lbcr.20190903.
-starting with PA=0 LBC-B 052616 LBC-R 052611

UT 05:27 - PA 180 LBC-B 052819 LBC-R 052816

UT 05:28 - PA 90 LBC-B 053001 LBC-R 052958 (weird quad streak through top half of both images - looks like laser beams but obviously its not)

UT 05:31 - PA 270 LBC-B 053211 LBC-R 053208

UT 05:37 - co-pointing script for Dumbell Nebula - will run dofpia here first

UT 05:41 - running dofpia

UT 05:45 - dofpia complained of no good pupils found due to big cloud passing through FOV

again - couldn't find data

Waiting for OB to finish and data files to appear in /newdata
Waiting for file in /newdata/lbc?.20190903.??????.fits
******* WARNING *******
Data files DID NOT appear in /newdata
-but data was in /newdata, pressed any key and it found it

UT 05:50 - cloud passed lots of pupils in collimation image
pressed shift+s in dofpia - closing telescope due to clouds

UT 05:51 - Used LBC.rotcenter task to compute rotator centers from 0 & 180 and 90 & 270 image pairs.

Blue rotator center is roughly (1037,2921.5) from 0/180 pair or (1038,2919.5) from 90/270 pair

Red RC is roughly (1081.5,2913) from 0/180 and (1081,2912) from 90/270 pair.

CRPIX1,2 in headers are (1035,2924) for Blue (1078,2913) for Red so the differences are small: 2-5 pix.

UT 06:59 - still closed

UT 07:11 - O. Kuhn running MODS cal tests (updating scripts and checking exposure times)

UT 07:12 - going to open again

-will try the Crescent Nebula RA 20 12 Dec +38 00

UT 07:20 - moving to co-pointing field for collimation

UT 07:22 - running DOFPIA
-LBC-B hard flushed twice and was slower in taking exposure

LBC-B 072258 - full Chip2 was read out, but LBC-R ran the strip mode of Chip2, following LBC-B was right size for collimation

-Red pupil looks triangular

UT 07:33 - DOFPIA still going

UT 07:35 - sending co-pointing script

UT 07:37 - co-pointing done

UT 07:38 - starting Crescent Nebula script

UT 07:40 - LBC-B is an exposure behind, really slow - not actually taking an exposure

several error messages on Blue: CCDCTRL bad input voltage line 7 (+23.90V instead of +25.0V+-1.0V)
  2019/09/03 07:41:05.114977     >>> preset/expose/save interrupted at expose time
UT 07:41 - LBC-B finally started an exposure

UT 07:44 - only took 1/2 Uspec exposures

UT 07:45 first dither on blue (2nd offset position)
  2019/09/03 07:45:37.880000 CAMERA CCDCTRL bad input voltage line 7 (+23.90V instead of +25.0V+-1.0V)

  2019/09/03 07:46:27.991480     >>> preset/expose/save interrupted at expose time
-LBC-B keeps hard flushing after every attempt, keeps trying to find pupils and increasing guider exposure time

UT 07:50 - LBC-B still hasn't taken an exposure at the 2nd dither position
96 2019/09/03 07:51:05.794833     camera system encountered a problem: CCD Controller error (9) [src/channel/channel.c:1247]
95 2019/09/03 07:51:05.794761 CAMERA   CCD Controller error (2) [src/camera/camera.c:1259]
94 2019/09/03 07:51:05.794719 CAMERA   reinitializing CCD Controller failed
-LBC-B tracker images are bad - shows electronic noise

UT 07:51 - closing for clouds

UT 08:02 - powered LBCs off, running lbckill/lbcstart, powering on again

UT 08:06 - Blue Channel Camera CCD controller error - failed to initialize:

-had a number of the following errors:
8 2019/09/03 08:06:10.646000 CAMERA CCDCTRL bad input voltage line 7 (+23.90V instead of +25.0V+-1.0V)
25 2019/09/03 08:06:28.285476     CameraInitialize failed, disabling CAMERA system
24 2019/09/03 08:06:28.285415     camera system encountered a problem: CCD Controller error (9) [src/channel/channel.c:527]
23 2019/09/03 08:06:21.394686 CAMERA   ccdctrlinit failed; CCD Controller error, (8) [src/camera/camera.c:560]
UT 08:08 - LBCs powered off, powering back on

UT 08:12 - LBCs came up OK - no errors

This is voltage line 7 normally when it comes up:
154 2019/09/03 08:11:35.411000 CAMERA CCDCTRL input voltage line 7 is +24.13V (+25.0V+-1.0V)
- may need to adjust power supply to fix this issue

-IT being generated

UT 08:20 - open again

UT 08:21- testing NS guiding - going to collimate in field near object RA 02:02 Dec +32:41

UT 08:24 - cleared out the DX +2000 z11 (cleared active optics, then changed global z11 to 0 and clicked update), also cleared active optics on SX

UT 08:25 - running dofpia

UT 08:29 - DX offset in the middle of DOFPIA

UT 08:31 - DX offset back again to original position

UT 08:32 Waiting for file in /newdata/lbc?.20190903.??????.fits
******* WARNING *******
Data files DID NOT appear in /newdata
-files were in newdata
UT 08:35 - DX moved again

UT 08:37 - running DOFPIA again, SX looks pretty bad

-blue thinks it converged but lots of astigmatism and coma

UT 08:41 - clearing active optics SX and running dohybrid

-Blue looks like its rotated the image 90 degrees - still fitting the bright part of the pupil as if it were an elongated pupil with astigmatism

Waiting for file in /newdata/lbc?.20190903.??????.fits
******* WARNING *******
Data files DID NOT appear in /newdata
- data are in /newdata

-switched to the FPIA portion of dohybrid and the blue side looks good (better than its been all night)

-then it got worse, then slightly better (but added -785 in z11)

-Shift+S in dohybrid to stop and proceed with NS testing

UT 08:54 - running co-pointing script

UT 08:58 - moving to NS target for C. Veillet 2018YH (hazardous rock recovery search)

-LBC-B - guiding is delaying same issue again (IT 7829)
36 2019/09/03 09:00:55.225000 CAMERA CCDCTRL bad input voltage line 7 (+23.90V instead of +25.0V+-1.0V)
35 2019/09/03 09:00:32.860044     >>> preset/expose/save interrupted at expose time
-no blue data is being taken, red NS is being taken

1135,2837 in image 1 (O. Kuhn spotted it) - same position on detector in image 2

UT 09:08 - LBC-B first image finally being taken (yay?)

UT 09:12 - lbcb.20190903.090656 - too few pixels downloaded - can see discontinuities in image (first LBC-B image) - chip 2 on LBC-R appears as Chip 1 in LBC-B, Chip 3 on LBC-R appears to be Chip 2 on LBC-B, Chip 1 on LBC-R appears as Chip 4 ob LBC-B

lbcb.2090903.091116 - looks OK
55 2019/09/03 09:17:12.380975     OB execution error: one or more systems failed [src/lbc/actions.c:1162]
54 2019/09/03 09:17:12.280041     >>> preset/expose/save interrupted at expose time
-OB stopped itself after this error - lets end this

UT 09:18 - reconfiguring for MODS-Bino

UT 09:19 - powering off LBCs

MODS-Bino

UT 09:35 - MODS configured putting into observing mode and homing AGWs

-going to check the hotspot center on AGWs

UT 09:40 - test preset

UT 09:42 - error on preset - no AGW unit selected (both MODS got this error) - restarting GCS

-guidecamera and wavefront sensor error

UT 09:44 - restart azcam server

UT 09:55 - logged into azcamserver via VNC and selected start azcam server icon on dekstop

UT 09:56 - MODS-1 AGW5 - 5w (wavefront sensor) shows "Error Controller reset error check power and fibers) - client connection closed

-rebooted mods-1cam computer

-stopping/starting GCS left

UT 10:03 - on azcamserver computer 5g/5w showed some connection, but then it stopped

-stopping/starting GCS left again

UT 10:07 - O. Kuhn found that the HEB on MODS-1 shut off, turning it back on, restarting rest of MODS services (not sure when/how they went off) - restarting GCS left again, 5w/5g still not working

-When did MODS-1 shutoff services? last exposure with MODS1 was UT 07:10

UT 10:22 - authorizing MODS-2 only

-pointing and collimation check

UT 10:34 - attempting to collilmate - too much coma

MODS2 hotspot/rotator center check

-xcenter 182 ycenter 163

[85:449,51:424] readout DETSEC
Result = 267,214

lbto@obs2:18 % gcsclient getHotspot right
Hotspot is (x,y): 267.00, 240.00
0.06414 mm/pix for MODS

off by 26 pixels = 1.66 mm

-->next step do an observation to see if we are really off by this amount

Going to move to StoneA - use mask

Computed Mask Alignment Offset:
   dX = 3.851 arcsec
   dY = -1.475 arcsec
  dPA = -0.018 deg

MODS2 Offset Command:
   offsetpointing -0.0183  3.851 -1.475 detxy rel

-need to try other astrometric mask to confirm rotator zp - TBD

UT 11:13 - going to do spectrophotometric GD71 - run as many modes as we have time to do

UT 11:28 - starting spectroscopic exposures

UT 11:42 - restarting for MODS2 (accidentally sent to MODS1)

-used observing script for g191b2b instead of GD71, so exposures times are off (as GD71 is 1.2 mags fainter) and flux is lower (but MODS2B is chugging along just fine!)

UT 12:03 - restarted script but properly named for GD71 - starting with red grating

UT 12:09 - aborting spec-std switching to do MODS2 flats (g, r)

-good set for both filters

UT 12:32- skyflats for u,z

-splotch in top right corner for all filters, dust ring on u only

-mods2r 0050 saturated

-mods2b collected u,g skyflats and mods2r collected r,z skyflats.

UT 12:47 - closing enclosure

-taking MODS2 biases 3kx3k


NOTE: Sky brightness monitor is stuck at 21.89

-- BarryRothberg - 02 Sep 2019
Topic revision: r12 - 03 Sep 2019, BarryRothberg
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback